A circular parallel plate capacitor has a capacitance of 144 pF and a separation distance of 1.50 mm. (a) What radius should this capacitor's plates have? (b) If a potential difference of 12.0 V is applied, what load does each board store? Ans: (a) r = 9.0 cm, (b) Q = 2.0 nC
Answer #1

a)

C = eo A / d

144* 10^-12 = 8.85* 10^-12* pi * r^2 / ( 1.5* 10^-3)

r = 0.088 m

r = 9 cm

======

b)

q = CV = 144* 12 = 1.7 * 10^-9

q = 2 nC

=======

Comment before rate in case any doubt, will reply for sure.. goodluck